# Student-AI Interaction: A Case Study of CS1 students

### v1

@@ 3.4. Study Procedure
We provided participants with three programming tasks in an office within the computer science department during the spring semester of 2024. The programming tasks were tailored to the complexity appropriate for a CS1 level and covered topics taught by the instructor throughout the semester preceding the study. Each participant had one hour to respond to the three programming tasks, utilizing the Python programming language that they had learned during the semester. Prior to commencing programming, we offered instructions on how to utilize the plug-in for inputting prompts. It was emphasized that no other Generative AI or aids were permitted during this time. These tasks were designed to encompass several fundamental programming concepts. The first task involved identifying and extracting even numbers from a given set, requiring an understanding of conditional statements and loops for iteration. The second task introduced input/output operations and string manipulation, challenging users to read input from the user and reverse the order of characters. Lastly, the third task delved into the concept of functions and sequences, prompting users to create a function that generates the first n values of the triangular numbers sequence. Each task touched upon recursion or iterative approaches, emphasizing algorithmic thinking and problem-solving skills. Collectively, these tasks addressed core programming concepts such as conditionals, loops, functions, input/output operations, and algorithm design, providing participants with a comprehensive understanding of essential programming principles.
@@
3.5. Data Instrumentation
The plug-in (Nam et al., 2024) we used for this study logs detailed user interactions within the text editor, including all changes to text selections and document modifications, along with the precise timestamps of these events. It also records the prompts users write and the responses generated by ChatGPT. The plug-in interface provides a prompt-writing box on the left side of VSCode, allowing users to access ChatGPT by writing prompts. Additionally, users can see Generative AI responses and their code in the same window, making it easy to compare the Generative AI responses with their code. This setup ensures a focused environment where users are less tempted to use other AIs. Figure 1 shows the interface of the plug-in.
@@ 3.6. Data Analysis
To address our research inquiries, we conducted both qualitative and quantitative analyses of the data. Two authors independently coded the data for three programming tasks, each varying in difficulty levels and topics, utilizing the initial codebook. Subsequently, they engaged in discussions to review the initial coding outcomes, resolve any conflicts, and enhance the codebook further. With the identification of each new code, we reviewed all prior comments for possible adjustments.
@@
We conducted two surveys, one administered before the programming session and another after, to assess participants’ level of self-efficacy. Participants rated their self-efficacy levels using Likert scales, indicating their agreement with statements related to their programming abilities. These Likert scores provided quantitative data to measure changes in participants’ self-efficacy before and after engaging in the programming tasks.
@@ 4. Results
Fifteen CS1 students participated in our students, and we asked each of them to do their best to complete three programming tasks. Participants P3, P5, P6, P9, P12, and P14 solved the three programming tasks successfully. All participants except P3 and P15 had at least one correct answer for the three tasks. Notably, participants P1, P4, P12, and P13 did not complete their last tasks, and P13 also did not complete their second task. Overall, we analyzed 40 solutions submitted by participants, from which 29 solutions were created with assistance from the Generative AI plug-in provided, and 11 without Generative AI.
4.1. Frequency and Adoption Time of Generative AI
We assigned one point to each user request to track how many times users wrote prompts using Generative AI. Table 2 displays the frequency of participants using the Generative AI through ChatGPT plug-in across the three programming tasks in our study, illustrating the different levels of dependence on Generative AI assistance among participants.
@@ 4.1. Frequency and Adoption Time of Generative AI
Table 2 shows P2 made more extensive use of Generative AI compared to other participants. P2 used ChatGPT nine times to solve Q1 but was unable to arrive at a correct solution. This instance shows that despite multiple interactions with Generative AI, the user was not able to solve the task. Participants’ strategies in using Generative AI and their inherent problem-solving abilities both play crucial roles in their success. Some participants, like P3 and P6, achieved correct answers (’C’) without relying on Generative AI. P3, for instance, solved all their tasks correctly without using Generative AI for two out of three programming tasks, indicating that some participants may not need Generative AI to perform well.
@@
As participants did not have any limitations on using Generative AI in terms of when and how many times it was used, we observed that users typically engaged with generative AI at three different stages.
@@
At the beginning of the programming. During the early stages of the programming, participants in 18 tasks turned to ChatGPT right away, without making any initial attempts on their own. Typically, they would type in the full description of the task or paraphrase it. P9, P10, P14, and P15 started using Generative AI right from the beginning, asking ChatGPT to help complete the entire task for them for all three programming tasks.
@@
Middle of programming. In this study, for n=9 questions, Generative AI users sought assistance of Generative AI in the middle of their programming. Initially coding manually, they resorted to using Generative AI due to encountering errors with invalid code or programming concepts.
@@
After solving the problem. Interestingly, two participants (P3 and P5) employed Generative AI even after they had already arrived at the correct answer on their own. P3, for example, used Generative AI to validate she understood the problem description correctly by inquiring about the distinction between ’i’ and ’n’ in the triangular numbers sequence. P5 reached the correct solution on their own first, but still sought Generative AI’s solution for the entire question, seemed that they attempted to verify the correctness of their solution by comparing their solution with the Generative AI’s responses. Table 3 shows when Generative AI users utilized Generative AI.
User When (Generative AI Used) Approach
Q1 Q2 Q3 Q1 Q2 Q3
p1 - Middle - - Hybrid/partial -
p2 Early Middle Early Step by step Hybrid/partial Full description
p3 - - After-Solving - - Step by step
p4 - - - - - -
p5 - After-Solving Early - Full description Step by step
p6 - Middle - Themselves Hybrid/partial -
p7 Middle Middle Middle Hybrid/partial Hybrid/partial Hybrid/partial
p8 - - Middle - - Hybrid/partial
p9 Early Early Early Full description Full description Full description
p10 Early Early Early Full description Full description Full description
p11 - Early Early - Full description Full description
p12 Early Middle - Full description Hybrid/partial -
p13 Middle - - Step by step - -
p14 Early Early Early Full description Full description Full description
p15 Early Early Early Full description Full description Full description
@@
Table 3. Summary of Generative AI Usage and Approaches by Participants: The table summarizes the Generative AI usage and approaches taken by users for different programming questions (Q1, Q2, Q3). The first set of columns under ”When (Generative AI Used)” indicates the timing of Generative AI usage classified as ”Early”, ”Middle”, ”After-Solving”, or ”-” if not used. The second set of columns under ”Approach” details the approach taken by participants, such as ”Hybrid/partial”, ”Step by step”, ”Full description”, and ”-” if not used.
4.2. Users Actions
We identified six common actions users performed when using Generative AI, namely, reading, thinking, writing code, modifying code, prompting, and debugging. Table 4 describes these actions.
Table 4. Description of six common actions users performed when using Generative AI user actions.
Action Description
@@ 4.2. Users Actions
Thinking Thinking about and confirming a suggestion
@@
The actions described in Table 4 varied among participants, with some repeating actions multiple times and others performing them only once. Figure 2 shows the actions of two participants. Participant P14 consistently followed a straight path across all three programming questions, performing each action once. In contrast, P7 in question 4 repeated certain actions such as running, debugging, prompting, modifying, and running again. Overall, we observed 17 instances of straight actions and 23 instances of repeated actions.
(a) Refer to caption
(b) Refer to caption
Figure 2. Actions of participants: (a) Linear Actions of P14 for all questions, (b) Non-linear Actions of P7 for Question 3
4.3. Approach in Interacting Generative AI
We categorize users’ approaches to Generative AI based on prompts. The table 3 shows how users approach Generative AI. We identified three strategies among users: full description, step-by-step, and partial. In the full description, the user copies the full description of the task into the plugin. In step-by-step, participants broke the question down into structured subgoals and used Generative AI to solve the problem step by step. We also observed a hybrid or partial approach, where participants tackled some parts of the question independently while seeking assistance from Generative AI for some other parts.
Full description
In total, we have n=17 approaches to Generative AI that utilized the entire of question in their prompt. P2 employed a full description of question 3, while for questions 1 and 2, their approach was step-by-step and Hybrid/partial. P5 only utilized the full description for question 2, although they already had the correct answer before resorting to Generative AI. P9, P10, P11, P14, and P15 initially utilized Generative AI by transcribing the full description of the question without writing any meaningful lines of code at the beginning of their programming process. Using a full description of the question does not necessarily guarantee that they will be able to find the correct answer. Among the participants who relied on full descriptions, P10 for Q3, P15 for Q1, Q2, and Q3 were unable to find the correct answers. The finding in Table 3 reveals that the full description approach is commonly used in early programming for most cases.
Step-by-step
@@ 4.3. Approach in Interacting Generative AI
Some users attempted to approach the questions step-by-step, breaking them down into various sub-goals. P2, P3, P5, and P13 adopted a similar Generative AI approach to some of the problems. For example, user P5 used Generative AI for the first time to inquire, ’How to round numbers in Python?’ and then followed up with a second query, “How to round numbers in Python only from the tenths place?”. Similarly, P13 used Generative AI three times for their first question. The first inquiry was, “In Python, can you turn an integer into a list?” The second query was about arrays: “What is an array?” Lastly, the user asked: “Given an integer n, produce the array: 0123…n, 1234…0, n012…n-1?.”
Hybrid/Partial
@@
Some participants attempted to write programs independently, but relied on Generative AI for debugging to resolve errors or others. We have identified n=8 approaches that we categorize as Hybrid or Partial. For example, User P2 asked Generative AI, “Fix my code to work.” User P6 encountered an error and copied the message: "TypeError: ’int’ object is not subscriptable".
4.4. Prompt Category
We analyzed the prompts that users submitted to ChatGPT each time they engaged with it. A total of 15 participants wrote 60 prompts. We categorized the prompts based on their relation to questions. The categories include Coding Concepts in Python, Program Logic, Debugging, and entire of Question.
Coding Concepts in Python.
In our study, nearly a third of the prompts (19 out of 60) were related to coding concepts. Seven users sought assistance from Generative AI specifically for coding concepts, indicating their understanding of logic but their need for coding support. Notable inquiries include P2’s query, “How to append to a list?”, P5’s question, “How to round numbers in Python?”, P6’s curiosity about “Python built-in function for arranging list in increasing order”, and P7 asking, “In my program, I am trying to move the negative numbers to the front. Could I initialize an empty list first?”
Program Logic.
@@ 4.4. Prompt Category
In 10 prompts, participants utilized Generative AI because they found it challenging to determine the sequence of steps or instructions needed to accomplish a specific task or goal within the program. For instance, P2 asked about ”a function to return the first n values of the triangular number sequence starting from 1.” P12 needed help trying to reverse a list. P7 inquired, ”I am trying to produce the array 0, 1, 2, 3 all the way to n, meaning it could also be n, 0, 1, 2. Is my approach correct so far?”
Debugging.
@@
Some novices turned to Generative AI for error resolution, indicating that their prompts typically revolved around the question: “How can I solve this error?” In total, 11 prompts were utilized Generative AI for error resolution. Often, users would simply copy the error message from the console and ask ChatGPT for help. P6 copied and pasted the code, asking, “What is wrong with this code?” P10 asked, “Why does the code not work?” P6 specifically inquired about the error received: “TypeError: ‘int’ object is not subscriptable.”
Entire of Problem description
Some Generative AI users simply copied the question descriptions, which we considered as their entire prompt in the question category. A total of 20 out of 60 prompts fell into this category. In some cases, such as P11, asking the full questions 2 and 3 resulted in receiving the entire answer with just one prompt.
4.5. Users’ reaction to the Generative AI responses
@@
We identify various user actions upon receiving responses from ChatGPT, which we categorize as acceptance categories. These actions reflect users’ reactions to Generative AI answers, and this collaboration can result in either successes or failures. These categories are based on the prompts and their reactions to the Generative AI’s answers.
4.5.1. Acceptance Category
Complete use
@@ 4.5.1. Acceptance Category
Occasionally, participants completely accept the Generative AI’s response in its entirety. We have $n=18$ prompts where users accept Generative AI answers based on these prompts. For example, P10 formulated a prompt for Question 3 “Given an integer n, produce the array” and accepted the Generative AI’s solution without any modification. Similarly, P7, when addressing the same question, queried “I am trying to produce the array 0,1,2,3 all the way to n, meaning it could also be n,0,1,2 is my correct so far?” fully embraced the Generative AI’s response.
Selective use
@@
In this scenario, some participants tried to understand the Generative AI response and figure out how to implement it in their own way. There are $n=23$ prompts where users get ideas from Generative AI responses to write their own code. For instance, P5 asked: “How to round numbers in Python only from the tenths place” and the Generative AI response was “To round numbers in Python only from the tenths place, you can use the round() function”, followed by an example of how round() function works. The user completed the task on their code.
Retry
@@
Using recorded videos of facial expressions and eye movements, we identified cases in which participants spent an extended amount of time reading and (most likely) contemplating Generative AI’s responses. We observed six participants visibly contemplating the Generative AI’s response to 19 prompts. These participants read the response but did not copy the code or write it in their own way. In such cases, users attempted to write another prompt, which may indicate their rejection of the Generative AI solution. P12’s initial query was “I need help trying to reverse a list.” After reading the Generative AI response, the user attempted to write another prompt, “Help me turn a sequence of numbers into a list.” Once again, the user read the Generative AI’s response. Subsequently, P12 revised his prompt and wrote: “Without using a built-in function, help me turn a sequence of numbers into a list.” After this prompt, he began to write code in their own way.
4.6. Correctness of Generative AI-assisted solution
To assess the frequency of successful collaboration facilitated by Generative AI, we used manual inspection and testing of the sample inputs to evaluate the correctness of the submissions. We considered all the cases that the participants used Generative AI at any stage in problem-solving. Interestingly, among the $n=29$ solutions in which participants used Generative AI, 19 solutions ($65\%$) were correct and the rest (35%) were incorrect. Most of the problems stem from incorrect copy-paste from the ChatGPT plugin, and the rest were from the difficulty in integrating portions of the ChatGPT responses into the participants’ current solution.
4.7. Participants’ Self-efficacy

### v2

@@ 3.2. Study Procedure
The study was conducted in an office within the computer science department during the spring semester of 2024, with one participant per study session. The moderator conducting the study was not part of the CS1 course in any capacity (instructor, tutor, or TA).
Each study session began with a study debriefing, and the participant signing the informed consent form. The participant then filled out a pre-study questionnaire, which contained questions about: 1) demographics, 2) programming background, and 3) self-efficacy questionnaire(from (Pintrich et al., 1991)).
Then, the participants were introduced to the study environment. We chose the Python programming language and VSCode environment for the study, since it was already used in the course. Participants were also introduced to the plug-in, and were briefed on how to ensert prompts into it, should they choose to use Generative AI for their tasks. We also emphasized to participants that other Generative AI or aids (e.g., web searches) were not allowed during this time. We also clarified that their performance would not be judged by others, or affect their course grades, to ensure that the participation was non-coercive.
Participants were introduced to the programming tasks on a sheet of paper, and were given one hour to complete the tasks. On completion of the tasks, the participants completed a post-study survey containing the same self-efficacy questionnaire administered as part of the pre-study survey. The study sessions lasted between 30 and 70 minutes each.
3.3. Programming tasks
Participants were asked to complete three programming tasks, each of which is tailored to the complexity appropriate for a CS1 level and covered topics taught by the instructor throughout the semester preceding the study.
Participants were provided with input and output examples for each task. Collectively, these three tasks addressed core programming concepts such as conditionals, loops, functions, input/output operations, and simple algorithm design, providing participants with a comprehensive understanding of essential programming principles and problem-solving approaches.
As mentioned earlier, participants had a choice of using Generative AI or not to complete these tasks. The use of Generative AI had to be via the custom plug-in we developed, and the use of any other resource (e.g., web search) was disallowed.
3.4. Generative AI plugin
The goal of this study is to understand when, how, and why CS1 students interact with Generative AI tools and how effective these interactions are. Therefore, we decided to allow only the use of Generative AI for students’ help seeking to complete their tasks and disallowed all other sources.
In deciding what Generative AI tool to use for the study, we had several choices: OpenAI ChatGPT, Google Gemini, and Github Copilot, to name a few. We chose to use OpenAI ChatGPT for two reasons: 1) it provided general help beyond programming, and 2) from a prior study, it seemed to be the most familiar Generative AI among students (Amoozadeh et al., 2024).
However, in a recent study, Tankelevitch et al. (Tankelevitch et al., 2024) highlighted the metacognitive demands placed on users when interacting with Generative AI systems (e.g., the need to remember sub-tasks and which one they are on), and suggested the need to minimize switching between the task and the Generative AI help-seeking contexts. Therefore, we used a GPT-3.5 plug-in within the VSCode [^1] https://code.visualstudio.com/ (Nam et al., 2024) programming environment, to reduce the cognitive load associated with switching between different environments, thereby allowing participants to remain within a single, cohesive coding environment.
Figure 1 shows the VSCode plugin  (Nam et al., 2024) we used for this study. As the figure shows, the plugin had a simple interface comprising of a prompt-writing box (See ”Ask a question…” in the image), allowing users to access ChatGPT within the VSCode IDE. This prompt box allows students to freely input prompts without any restrictions on the number or type of questions, or prompt lengths. The Generative AI responses to the prompts are also provided in the same pane, as the figure shows. The fact that users can access Generative AI and see its responses together with their code in the same window allows for a focused environment where users do not need to switch context to another window, and are less tempted to use other AIs.
@@ 3.4. Generative AI plugin
3.5. Data gathering
Our study employed a multi-faceted approach to data collection, designed to capture a comprehensive view of participants’ interactions with Generative AI (here, ChatGPT) during programming tasks. We utilized a custom VSCode plugin to log all text editor interactions (including any text selections) and ChatGPT prompts, and responses. The audio and video recordings of the participants, along with recordings of the participants’ screens throughout the session provided qualitative data on problem-solving behaviors and reactions to AI assistance. Pre-study and post-study surveys collected demographic information and self-efficacy data using Likert scales. In addition, we collected detailed logs of ChatGPT conversations, including the full text of prompts and responses. This diverse dataset allowed us to analyze the frequency and timing of Generative AI use (RQ1), students’ activities and their patterns (RQ2), their Generative AI use strategies (RQ3), and the effects of Generative AI use on students’ self-efficacy (RQ4). By combining automated logging, video analysis, and self-reported measures, we ensured a foundation for both quantitative and qualitative analysis of novice programmers’ engagement with AI-assisted coding.
@@ 3.6. Data analysis
To answer our research questions, we conducted both qualitative and quantitative analyses of the data.
@@
Two authors independently coded the video data for three programming tasks, each varying in difficulty levels and topics, utilizing the initial codebook. Subsequently, they engaged in discussions to review the initial coding outcomes, resolve any conflicts, and enhance the codebook further. With the identification of each new code, we reviewed all prior comments for possible adjustments.
@@
We conducted two surveys, one administered before the programming session and another after, to assess participants’ level of self-efficacy. Participants rated their self-efficacy levels using 5-level Likert scales, indicating their agreement with statements related to their programming abilities. These Likert scores provided quantitative data to measure changes in participants’ self-efficacy before and after engaging in the programming tasks.
@@ 4. Results
Fifteen CS1 students participated in our study, and we asked each of them to do their best to complete three programming tasks. Table 2 summarizes the tasks completed and the correctness of solutions for each participant. Notably, most participants (P3, P15) had atleast one correct solution, but only 6 out of 15 participants (P3, P5, P6, P9, P12, P14) solved the three programming tasks successfully. Overall, we had 40 completed (correct + incorrect) solutions submitted by participants. We then analyzed the logs from the Generative AI.
4.1. RQ1: How frequently do students use Generative AI while programming?
We used the logs from the plugin to analyze the frequency of Generative AI use. Of the 40 completed participant submissions, 29 solutions were created with assistance from the Generative AI plug-in provided, and 11 without Generative AI. requests, as a measure of the frequency of participants’ Generative AI use during each task. Table 2 displays this frequency of participants’ plugin-aided Generative AI use, illustrating the different levels of dependence on Generative AI assistance among participants.
@@ 4.1. RQ1: How frequently do students use Generative AI while programming?
However, these use frequencies did not always translate into task completion successes. Overall, among the 29 Generative AI assisted solutions created by participants, only 19 solutions ($65\%$) were correct and the rest (35%) were incorrect. Even among two participants (P2, P5) used Generative AI extensively for a task, only one instance (P5, Q3) was successful; P2 did not manage to complete Q1 correctly, even after 9 Generative AI queries. Moreover, some participants, such as P3 and P6, achieved correct answers (’C’) for one or all tasks even without relying on Generative AI, indicating that some participants may not need Generative AI to perform well.
Thus, participants’ frequency of Generative AI use alone did not lead to successes; instead, participants’ strategies in using Generative AI and their inherent problem-solving abilities both played crucial roles in their success.
4.2. RQ2: How do students interact with Generative AI while programming?
To understand how and why some participants were able to use Generative AI to correctly complete tasks, and others were unsuccessful in doing so, we delved into various qualitative aspects of participants Generative AI use–when and how participants interacted with Generative AI, and how these translated into task completion successes.
4.2.1. When do students use Generative AI?
@@
We observed that participants typically engaged with generative AI at three different stages during their task:
@@
  • Early in the task. In our study, 8 participants, in a total of 18 tasks, turned to ChatGPT right away, at the beginning of the task, without making any initial attempts on their own; see instances of ”Early” in Table 3 under ”When”. Of them, four participants (P9, P10, P14, P15) used Generative AI right from the beginning for all three programming tasks. Typically, during these instances, participants typed in the description of the task as is, and relied on Generative AI to complete the task for them.
@@
  • Middle of the task. In 9 out of 45 total tasks, participants (7 out of 15) sought the assistance of Generative AI in the middle of their programming; these instances are labeled ”Middle” in Table 3, under ”When”. Initially coding manually, they resorted to using Generative AI when they encountered errors, or to understand programming concepts.
@@
  • After completing the task. Interestingly, two participants (P3 and P5) employed Generative AI even after they had already arrived at the correct answer on their own. See instances of ”After-Solving” in Table 3; Section ”When”. Here, P3 used Generative AI to validate she understood the problem description correctly by inquiring about the distinction between the terms (i, N) mentioned in the problem. P5 reached the correct solution on their own first, but still sought Generative AI’s solution for the entire question, apparently to verify the correctness and/or effectiveness of their solution by comparing their solution with the Generative AI’s responses.
User When (Generative AI Used) Approach Usage Pattern
Q1 Q2 Q3 Q1 Q2 Q3 Q1 Q2 Q3
P1 - Middle - - Hybrid/partial - I I -
P2 Early Middle Early step-by-step Hybrid/partial Full description I I I
P3 - - After-Solving - - step-by-step I I I
P4 - - - - - - I I -
P5 - After-Solving Early - Full description step-by-step I L I
P6 - Middle - Themselves Hybrid/partial - L I I
P7 Middle Middle Middle Hybrid/partial Hybrid/partial Hybrid/partial I L I
P8 - - Middle - - Hybrid/partial I I I
P9 Early Early Early Full description Full description Full description L L L
P10 Early Early Early Full description Full description Full description L I L
P11 - Early Early - Full description Full description L L L
P12 Early Middle - Full description Hybrid/partial - I I -
P13 Middle - - step-by-step - - I I -
P14 Early Early Early Full description Full description Full description L L L
P15 Early Early Early Full description Full description Full description L L L
@@
Table 3. Summary of Generative AI Usage and Approaches by Participants: The table summarizes the Generative AI usage and approaches taken by users for different programming questions (Q1, Q2, Q3). The first set of columns under ”When (Generative AI Used)” indicates the timing of Generative AI usage classified as ”Early”, ”Middle”, ”After-Solving”, or ”-” if not used. The second set of columns under ”Approach” details the approach taken by participants, such as ”Hybrid/partial”, ”Step by step”, ”Full description”, and ”-” if not used. The third column shows nature of participants’ activities, with some repeating activities multiple times as iterative ”I” and others performing them only once as linear shown as ”L”.
4.2.2. What are common Generative AI usage patterns among students?
We conducted an activity analysis of the students’ programming tasks, coding the activities they engaged in. We identified six common activities users performed when using Generative AI, namely, reading, thinking, writing code, modifying code, prompting, and debugging. Table 4 defines these activities. Participants engaged in these activities in varying levels, with some engaging in the same activity on multiple instances during the same task, and others performing them only once per task.
An analysis of the activity sequences for each task revealed two large patterns, as illustrated in 2; the figure shows the activity sequences [^2] Notice that in the figure ”correct” and ”wrong” are not activities; we also did not count ”running code” as an explicit activity because it was a part of task completion or verification. of two tasks by P14 and P7 respectively. Whereas P14 followed a straight path across all three programming questions, performing each activity once, P7 in question 4 repeated certain activities such as running, debugging, prompting, modifying, and running again. Overall, the 15 participants in our study completed a total of 40 tasks (Table 2, column Completed); of them, we observed 17 instances of straight or linear activity sequences and 23 instances of repetitive activities. These are indicated by “L” and “I” in Table 3, under ”Usage Pattern”.
Activity Description
@@ 4.2.2. What are common Generative AI usage patterns among students?
Thinking Thinking about solution, or Generative AI prompt and response
@@
Table 4. Description of six common activities users performed when using Generative AI user activities.
The pattern of a student returning to the same activity is an indicator of an iterative process (e.g., several iterations of re-running code, or prompting Generative AI more than once for the same task). In contrast, a linear pattern is an indicator of one-shot task completion, assisted with or without Generative AI. The fact that over 40% instances were linear is suggestive of one of two possibilities. The first is that of high task performance abilities, as marked by minimal help-seeking, and lack of edit-verify loops. The second possibility is that of over reliance on Generative AI, wherein participants made a single prompt to Generative AI the result of which resulted in a direct solution to the problem, following which participants did not have to enter an edit-verify loop. To assess which of these possibilities caused the observed usage patterns, we drilled down into the nature of participants’ interactions with Generative AI.
(a) Linear activities of P14 for all questions
(b) Iterative activities of P7 for Question 3
Figure 2. Examples of activities
4.2.3. What interaction strategies do students adopt with Generative AI?
There were three key aspects to participants interactions with Generative AI: the level of task decomposition, the kind of information they needed, and how participants exploited the Generative AI response.
Levels of task decomposition
As Table 3 shows, participants engaged in three prompting strategies, based on the level of task decomposition: full description, step-by-step, and partial:
  • In the full description strategy, the participant copied the full description of the task into the plugin, thereby offloading task completion entirely to Generative AI. In total, there were 17 instances across 9 participants where participants, early in the task (Table 3), provided the entire problem description to Generative AI to be solved by the latter. Of them, 5 participants adopted this approach for only one or two tasks, and relied on step-by-step or hybrid /partial approach for others. However, 4 participants (P9, P10, P11, P14, P15) entirely relied on this strategy for all tasks, taking to Generative AI very early in the task. Note, however, that this opportunistic approach to offload the task completion to Generative AI did not translate to complete and correct solutions. Among the participants who relied on full descriptions, P10 (for Q3) and P15 (for Q1, Q2, Q3) were unable to find the correct answers.
@@ 4.2.3. What interaction strategies do students adopt with Generative AI?
  • In the step-by-step interaction strategy, participants broke the question down into structured sub-goals and then used the Generative AI to solve the problem step by step–resulting in an iterative approach. Four participants (P2, P3, P5, P13) adopted this problem-solving strategy. For example, participant P5 used Generative AI for the first time to inquire, ’How to round numbers in Python?’ and then followed up with a second query, “How to round numbers in Python only from the tenths place?”. Similarly, P13 used Generative AI three times for their first question. The first inquiry was, “In Python, can you turn an integer into a list?” The second query was about arrays: “What is an array?” Lastly, the user asked: “Given an integer n, produce the array: 0123…n, 1234…0, n012…n-1?.” Indeed, there is diversity in the nature of prompts, as we describe later.
@@
  • The remaining participants used a hybrid or partial strategy where participants tackled some parts of the task independently while seeking assistance from Generative AI for some other parts. Notably, participants attempted to write programs independently, but turned to Generative AI for debugging to resolve errors or others. We identified n=8 instances that we categorized as Hybrid or Partial. For example, when stuck with a bug, participant P2 asked Generative AI, “Fix my code to work.” User P6 encountered an error and copied the message: "TypeError: ’int’ object is not subscriptable". We observed only one instance of a participant successfully completing a task on their own (Table 3, Q1, P6 labeled ”Themselves”).
Information needs.
We analyzed the prompts that users submitted to ChatGPT, via the logger. In all, 15 participants wrote 60 prompts, and we categorized them based on the information need the prompt aimed to serve.
  • Entire Solution. As described earlier, several participants simply typed in the question descriptions, which we considered as their entire prompt. These accounted for a third of all prompts (20 out of 60). Often, these prompts were issued early on during the task, and participants obtained the entire solution to the task, via this single prompt; 8 out of 15 participants engaged in this behavior.
  • Coding Concepts. Another popular prompting strategy, namely seeking to understanding programming concepts, was equally common among participants. About a half of participants (7 out of 15) accounting sought such information, accounting for  30% (19 out of 60) of all prompts. These instances of conceptual understanding, arose in largely two cases, namely when the participant had a logical understanding of what to do, but needed help executing them in Python (e.g., P2: “How to append to a list?”, P5:“How to round numbers in Python?”), to understand whether the language allowed something (e.g., P7: “In my program, I am trying to move the negative numbers to the front. Could I initialize an empty list first?”), or to understand jargon (e.g., P13: “What is an array?’).
@@
  • Program Logic. In about 15% of the time (10 out of 60 prompts), participants utilized Generative AI because they found it challenging to determine the sequence of steps or instructions needed to accomplish a specific task or goal within the program. In other words, they used Generative AI for planning help. For instance, P2 asked about ”a function to return the first n values of the triangular number sequence starting from 1.”, expecting to receive the steps to accomplish the task. P12 needed help trying to reverse a list. P7 inquired, ”I am trying to produce the array 0, 1, 2, 3 all the way to n, meaning it could also be n, 0, 1, 2. Is my approach correct so far?”.
@@
  • Debugging. Finally, some participants turned to Generative AI for debugging and their questions typically revolved around the question: “How can I solve this error?”. In total, 11 out of 60 prompts utilized Generative AI for error resolution. Often, users would simply copy the error message from the console and ask ChatGPT for help. P6 copied and pasted the code, asking, “What is wrong with this code?” P10 asked, “Why does the code not work?” P6 specifically inquired about the error received, in some ways treating it like a search engine: ‘‘TypeError: ‘int’ object is not subscriptable.’’
Exploiting Generative AI responses
@@
We identified various user actions upon receiving responses from Generative AI, which we categorize as acceptance categories. These actions reflect users’ reactions to Generative AI answers, and this collaboration can result in either successes or failures in the completion of tasks. These categories are based on the prompts and their reactions to the Generative AI’s answers.
@@
  • Entire response. Participants often accepted the Generative AI’s response in its entirety. We had $n=18$ out of 60 prompts we found that users accepted Generative AI answers, without any evidence of explicitly evaluating the responses. For example, P10 formulated a prompt for Question 3 “Given an integer n, produce the array” and accepted the Generative AI’s solution without any modification. Similarly, P7, when addressing the same question, queried “I am trying to produce the array 0,1,2,3 all the way to n, meaning it could also be n,0,1,2 is my correct so far?” fully embraced the Generative AI’s response.
@@
  • Selective use. Another common use of Generative AI responses was that participants tried to understand the Generative AI response, and then translate that understanding to implementing their own solution. There were $n=23$ out of 60 prompts where users sought ideas from Generative AI responses to write their own code. For instance, P5 asked: “How to round numbers in Python only from the tenths place” and the Generative AI response was “To round numbers in Python only from the tenths place, you can use the round() function”, followed by an example of how round() function works. The user then completed the task using their own code.
@@
  • Reject and Retry. Using recorded videos of facial expressions and eye movements, we also identified cases in which participants spent an extended amount of time reading and (most likely) contemplating Generative AI’s responses. We observed six participants visibly contemplating the Generative AI’s response to 19 (out of 60) prompts. These participants read the responses but did not exploit it, by way of either copying the code from the response, or simply write it in their own way. Instead, they went on to write another prompt, which may indicate their rejection of the Generative AI solution The reasons for this varied from Generative AI responses not meeting their information needs, to a rejection of the implementation choices made in the responses. P12’s initial query was “I need help trying to reverse a list.” After reading the Generative AI response, the user attempted to write another prompt, “Help me turn a sequence of numbers into a list.” Once again, the user read the Generative AI’s response. Subsequently, P12 revised his prompt and wrote: “Without using a built-in function, help me turn a sequence of numbers into a list.” After this prompt, he began to write code in their own way.
4.3. RQ3: How does student self-efficacy change before and after programming with Generative AI?
With such diversity in Generative AI usage among participants–in terms of frequencies, usage patterns and strategies, and success rates– we went beyond simple task completion to metacognition, specifically to evaluate the impact of Generative AI tools on students’ self-efficacy.
For this, we analyzed the self-efficacy questionnaires we administered as part of the pre-study and post-study surveys. Unfortunately, for the first five participants (P1 to P5), we did not capture the pre-study self-efficacy data; thus, we had data from 10 participants.